22) A firm has an effective (after-tax) cost of debt of 3%, and its weight of debt is 40%. Its equity cost of capital is 11%, and its weight of equity is 60%. Calculate the firm’s weighted average cost of capital (WACC). [Enter your answer as a decimal rounded to four decimal places.]

23) A firm is considering an investment project that costs $250,000 today and $250,000 in one year, but would produce benefits of $50,000 a year, starting in one year, forever. What is the NPV of this investment project if the firm applies an annual discount rate of 8% to all future cash flows?
Answer #1

22)

WACC = Weight of debt*after tax cost of debt + weight of equity*cost of equity

WACC = 0.4*0.03 + 0.6*0.11

WACC = 0.012 + 0.066

WACC = 0.0780

23)

Present value of cash inflow = Perpetual cash flow / discount rate

Present value of cash inflow = 50,000 / 0.08

Present value of cash inflow = 625,000

Present value of year 1 cash outflow = 250,000 / (1 + 0.08)

Present value of year 1 cash outflow = 231,481.4815

Total cash outflow = 250,000 + 231,481.4815 = 481,481.4815

NPV = Present value of cash inflows present value of cash outflows

NPV = 625,000 - 481,481.4815

NPV = $143,518.52